# A Comprehensive Survey of Direct Preference Optimization: Datasets, Theories, Variants, and Applications

### v2

@@ 5.1.1 Application on reasoning
Step-DPO (Lai et al., 2024) addresses this by introducing fine-grained supervision during fine-tuning. Instead of evaluating entire answers, Step-DPO treats individual reasoning steps as the units for optimization. This approach increases accuracy by focusing on identifying and correcting errors that typically arise mid-way through reasoning chains. Step-DPO maximizes the probability of selecting the correct next reasoning step while minimizing the probability of selecting incorrect steps, improving overall reasoning accuracy and factuality.
Step-Controlled DPO (Lu et al., 2024b) also enhances models’ reasoning ability in step-wise granularity. To be specific, Step-Controlled DPO generates step-wise errors(dispreferred sample) in the concrete step among the long reasoning steps. By deploying DPO in the training stage, SCDPO enables model to understand the specific step that starts to be erroneous, supporting the model to generate right response.
@@
DPO-augmented Self-Training(DPO-ST) (Wang et al., 2024c) strengthens the original self-training algorithm to enhance models’ ability of CoT reasoning. Self-training is based on the prospective that models can learn from their own outputs and enhance predictive accuracy or decision-making capabilities. Incorporating direct preference optimization into traditional self-training pipeline can improve models’ chain-of-thought reasoning ability.
Large language models exhibit varying capabilities in mathematical reasoning between dominant languages, such as English, and non-dominant languages. This disparity can largely be attributed to the imbalance in pre-training and fine-tuning data across different languages  (Chen et al., 2023b). To address this imbalance in reasoning abilities, MAPO  (She et al., 2024) introduces an innovative alignment method that leverages the robust reasoning capabilities of dominant languages to enhance the reasoning skills of non-dominant languages. By eliminating this requirement,  She et al. (2024) effectively streamline the reasoning process, making it more accessible. By utilizing both DPO and iterative DPO, MAPO significantly enhances multilingual reasoning capabilities, ensuring that models are not only more effective in their performance across languages but also more equitable in their reasoning abilities. This advancement paves the way for more inclusive applications of language models in diverse linguistic contexts.
@@ 5.1.2 Application on instruction-following
The instruction-following ability of large language model refers to their capacity to understand, interpret, and execute commands given to them in natural language (Lou et al., 2024b; Ouyang et al., 2022b). To state-of-the-art large language models, the instruction-following capacity is essential as LLMs cannot leverage their enormous knowledge to perform various tasks and interpret users’ meaning without instruction-following fine-tuning stage.The instruction-following ability of large language models (LLMs) is crucial for effectively executing commands in natural language. To enhance this capability, the AUTOIF (Dong et al., 2024) framework introduces two approaches: SFT + Offline DPO and SFT + Iterative Online DPO. For SFT + Offline DPO, after supervised fine-tuning, AUTOIF generates and filters various data scales, creating numerous positive and negative sample pairs for training. The AUTOIF framework employs both offline and iterative online DPO methods, significantly improving a model’s ability to understand and execute natural language instructions.  Yang et al. (2024b) propose an innovative data augmentation technique that simplifies complex instructions by breaking them down into manageable sub-components. These components are then modified and reassembled, preserving the original context and complexity while adding variability essential for training and evaluating the instruction-following capabilities of large language models. Their method can indeed improve the instruction-following capacity of large language models by generating nuanced instruction variants while maintaining the original topic and context, which also allows models to better understand complex instructions. DPO enables LLMs to better finish the decomposed instructions with human preference.
@@
LLMs struggle to generate responses that are aligned with given instructions, especially for complex instructions with multiple restraints. To handle the instruction following needs, Conifer (Sun et al., 2024) introduces a new dataset dedicated to improving the instruction-following ability under multiple complex restraints that are closed to the real-world application situations. Conifer constructs a brand-new dataset with GPT-4’s insights on instructions with complex restraints, making it accessible for instruction-following training. The models are further trained by direct preference optimization(DPO). Conifer-7B-DPO is aligned with DPO algorithm, achieving the goal of enhancing instruction following ability of LLMs.
@@
Large language models are prompt-sensitive as the quality of prompts greatly influences the instruction-following ability for LLMs’ performance on downstream tasks. The naive prompts induce the incorrect understanding of instructions. The free-form Free-form Instruction-oriented Prompt Optimization (FIPO; (Lu et al., 2024c)) enables models to follow the core instruction within the naive prompts that are considered as model-agnostic manner by accurately optimization. DPO is emplyed in the process of prompt optimization based on the large-scale prompt preference dataset, achieving remarkable instruction-following ability.
@@
With instruction-tuning, large language models learn to follow instructions to finish specific tasks. The ability or performance of LLMs can be evaluated on instruction following and corresponding responses. RS-DPO (Khaki et al., 2024) further imroves models’ instruction-following ability by combining rejection sampling and direct preference optimization. Evaluation on instruction-following benchmark demonstrates the effectiveness of thei combination in enhancing models’ performance.
@@ 5.1.3 Application on anti-hallucination
The conventional alignment fails to enhance LLMs’ factual accuracy and leads to the generation of false fact (i.e. hallucination) (Lin et al., 2024c). Moreover, former study demonstrated that introducing new knowledge or unfamiliar texts in fine-tuning stage can encourage large language models to hallucinate (Gekhman et al., 2024; Lin et al., 2024c), which makes SFT less factual as it trains on human labeled data that may be novel to LLM. Reinforcement learning also encourages LLM to hallucinate as the reward functions prefer longer responses on a diverse set of instructions that contain more details, which tends to stimulate the LLM to yield more false claims.
To tackle these factuality issues, alongside Retrieval-Augmented Generation (RAG; (Gao et al., 2023; Zhao et al., 2024a)),  Lin et al. (2024c) propose factuality-aware alignment (FLAME), which incorporates factuality-aware supervised fine-tuning (SFT) and reinforcement learning through direct preference optimization (DPO). DPO is central to this process, ensuring that LLM outputs align with factuality-aware objectives and enhance overall response quality. DPO, combined with factuality-aware alignment, offers a robust framework for improving the quality and factuality of LLM outputs.
@@ 5.1.4 Application on code-generation
Code language models(CLMs) are commonly evaluated by the passing rate to certain task or the accuracy to the generated code task. The running time for CLMs is not a standard criteria for code generation task. Code-Optimise (Gee et al., 2024) evaluates the quality of generated by considering both correctness and runtime. Code-Optimise is optimized through direct preference optimization fine-tuning by learning from two dimensional signals including preference pairs over Quick versus Slow as well as Passed versus Failed. An improvement for performance can be observed by DPO fine-tuning.
@@
Stable Code (Pinnaparaju et al., 2024) is a serials of based code models that is aimed at code understanding, code interpretation, code completion, mathematical reasoning, and relevant software engineering tasks. Stable Code’s programming capability is further reinforced through SFT and DPO. Training on the preference datasets such as UltraFeedback and HH-Anthropic dataset, Stable Code with 3B parameters outperform large language code models with 7B and 15B parameters, proving the effectiveness of DPO fine-tuning in post-training stage.
@@
Code LLMs usually generate executable code responses to finish the required engineering tasks, assisting human engineers to complete coding domain instructions. However, code LLMs are struggling to generate fast codes for real-world performance because code LLMs only treat code snippets as texts.  Nichols et al. (2024) introduce a new alignment to enhance code LLMs’ performance for generating fast code. Using the variant of DPO fine-tuning, Direct Performance Alignment (DPA; (Nichols et al., 2024)) fine-tunes code LLMs by training on faster and slower code paires for preference optimization, achieving new benchmark for code performance.
@@
Large language models have demonstrated comprehensive ability for various downstream tasks including code generation. General LLMs fails to accomplish all code generation tasks especially on competition level programming tasks as these tasks include many corner cases that are not trained in the pre-training stage. Instruct-Code-Llama (Liu et al., 2024e) introduces the new fine-tuning approach using Reinforcement Learning with Online Judging Feedback. In RL stage, Instruction-Code-Llama is further aligned using DPO with high-quality data, which achieves less complexity in time and space of generated codes and strengthens the efficiency. The above subsections elaborate the current application of DPO fine-tuning in downstream tasks for forefront LLMs research. Actually, DPO is also employed in evaluation task. RewardBench (Lambert et al., 2024) constructs a new benchmark for reward models to evaluate reward models’ capabilities in multiple domains such as chat, reasoning, safety as well as helpfulness and harmlessness. Direct judgement preference optimization (Wang et al., 2024e) deploys the DPO fine-tuning to enable the model to learn from the positve and negative samples, enhancing models’ evaluation capabilities across different domain of tasks including reasoning, judgement, and models’ response-comprehension.
@@ 5.2.1 Multi-modal Understanding
Image understanding: A prevalent issue in Large Vision Language Models (LVLMs) is hallucination, where the generated responses fail to align with the provided image and text contexts, significantly restricting the usages of LVLMs (Liu et al., 2024f; Yin et al., 2023; Cui et al., 2023b). Generally, humans prefer non-hallucinatory responses over hallucinatory ones. Consequently, some studies have reinterpreted the problem of model hallucinations as a preference optimization task, where DPO is widely employed.
@@
Yu et al. (2024a, b) proposed a token-level variant of DPO to mitigate hallucinations in dense captioning tasks. They found that some key words in the chosen responses required more intensive reinforcement. Specifically, they gathered human preferences  (Yu et al., 2024a) or AI feedback  (Yu et al., 2024b) in the form of segment-level corrections for hallucinations and implemented a dense variant of DPO.  Xiao et al. (2024) proposed a detect-then-rewrite pipeline for differentiating and mitigating hallucinations, and employed an variant of DPO which adaptively assigns each rejected response a hallucination severity score.
In Llama 3 Herd of Models (AI@Meta, 2024), authors adopted DPO in post-training stage and have find that instead of always freezing the reference model, updating it in an exponential moving average (EMA) fashion every k-steps helps the model learn more from the data, resulting in better performance in human evaluations. Besides, they observed that the vision DPO model consistently performs better than its SFT starting point in human evaluations for every finetuning iteration.
Video understanding:  Zhang et al. (2024b) introduced a novel framework named LLaVA-Hound that utilizes detailed video captions as a proxy of video content. Specifically, they extracted dense captions on video and then employed a large language model with these captions to score video QA. Their empirical results demonstrate this method’s robust alignment with the reward from GPT-4V (OpenAI, 2023a), which can take images (i.e. video frames) as input. Additionally,  Zhang et al. (2024d) developed LLaVA-Next-Video by supervised fine-tuning (SFT) LLaVA-Next-Image on video data, achieves better video understanding capabilities compared to LLaVA-Next-Image. Further, they employed LLaVA-Hound (Zhang et al., 2024b) which aligns the model response with AI feedback using DPO, showing significant performance boost.
Audio understanding: In Qwen2-Audio’s development (Chu et al., 2024), the authors directly employed vanilla DPO to align the SFT model with factuality and human preference.  Ye et al. (2024) developed a MLLM named CAT that responds to audio-visual content, and they have found that although MLLMs are equipped with multimodal understanding, they cannot perceive ambiguity. Therefore, they have reinterpreted ambiguity mitigation as a model preference optimization task and proposed an AI-assisted Ambiguity-aware Direct Preference Optimization (ADPO) method. Specifically, they took ambiguous responses that express the lack of clarity of specific audio-visual objects as negative responses, and then they utilized GPT to rewrite them into positive responses. After the multimodal training, they performed ADPO to align the model to generate the positive response (i.e. the precise description after the rewrite) and reject the negative response (i.e. the ambiguous description).
@@ 5.2.2 Multi-modal Generation
Previous works, such as  (Wallace et al., 2023; Yang et al., 2024c), have explored the align T2I generation using DPO. These approaches typically works under the bandit assumption, focusing on a latent reward applied to the entire diffusion denoise process. However, this ignores the sequential nature of the generation process. To address these limitations,  Yang et al. (2024e) propose a fine-grained approach with a dense reward. They develop a tractable objective focusing on the initial stages of the denoising chain. Besides, their method incorporates temporal discounting into implicit reward objectives, which better conform to the hierarchical structure of T2I generation.
Audio generation:  Majumder et al. (2024) leveraged an existing text-to-audio model, Tango, to synthetically create a preference dataset. In this dataset, each prompt is associated with a chosen audio output and several rejected audio outputs, which the diffusion model learns from. The rejected outputs theoretically contain some concepts from the prompt either missing or ordered incorrectly. The authors fine-tuned the text-to-audio model Tango with diffusion-DPO (Wallace et al., 2023) on the preference dataset. Their results demonstrate that this approach yields improved audio outputs over Tango, as evidenced by both automatic and manual evaluation metrics.
Current emotional text-to-speech (TTS) models  (Diatlova and Shutov, 2023; Li et al., 2024b) only learn the correct emotional outputs without fully comprehending other emotion characteristics, limiting their ability to capture the nuances between different emotions.  Gao et al. (2024) propose a controllable Emo-DPO approach, which employs direct preference optimization to differentiate subtle emotional nuances between emotions through optimizing towards preferred emotions over less preferred emotional ones.  Zhang et al. (2024c) have employed DPO in speech generation to align speech generation to human preferences, and they have found that DPO contributes to improvement in speech generation and validated that iterative DPO can consistently enhance the generated speech’s quality.

### v3

@@ 5.1.1 Application on reasoning
Step-DPO Lai et al. (2024) addresses this by introducing fine-grained supervision during fine-tuning. Instead of evaluating entire answers, Step-DPO treats individual reasoning steps as the units for optimization. This approach increases accuracy by focusing on identifying and correcting errors that typically arise mid-way through reasoning chains. Step-DPO maximizes the probability of selecting the correct next reasoning step while minimizing the probability of selecting incorrect steps, improving overall reasoning accuracy and factuality.
Step-Controlled DPO Lu et al. (2024b) also enhances models’ reasoning ability in step-wise granularity. To be specific, Step-Controlled DPO generates step-wise errors(dispreferred sample) in the concrete step among the long reasoning steps. By deploying DPO in the training stage, SCDPO enables model to understand the specific step that starts to be erroneous, supporting the model to generate right response.
@@
DPO-augmented Self-Training(DPO-ST) Wang et al. (2024c) strengthens the original self-training algorithm to enhance models’ ability of CoT reasoning. Self-training is based on the prospective that models can learn from their own outputs and enhance predictive accuracy or decision-making capabilities. Incorporating direct preference optimization into traditional self-training pipeline can improve models’ chain-of-thought reasoning ability.
Large language models exhibit varying capabilities in mathematical reasoning between dominant languages, such as English, and non-dominant languages. This disparity can largely be attributed to the imbalance in pre-training and fine-tuning data across different languages  Chen et al. (2023b). To address this imbalance in reasoning abilities, MAPO  She et al. (2024) introduces an innovative alignment method that leverages the robust reasoning capabilities of dominant languages to enhance the reasoning skills of non-dominant languages. By eliminating this requirement,  She et al. (2024) effectively streamline the reasoning process, making it more accessible. By utilizing both DPO and iterative DPO, MAPO significantly enhances multilingual reasoning capabilities, ensuring that models are not only more effective in their performance across languages but also more equitable in their reasoning abilities. This advancement paves the way for more inclusive applications of language models in diverse linguistic contexts.
@@ 5.1.2 Application on instruction-following
The instruction-following ability of large language models refers to their capacity to understand, interpret, and execute commands given to them in natural language Lou et al. (2024b); Ouyang et al. (2022b). To state-of-the-art large language models, the instruction-following capacity is essential as LLMs cannot leverage their enormous knowledge to perform various tasks and interpret users’ meaning without instruction-following fine-tuning stage.The instruction-following ability of LLMs is crucial for effectively executing commands in natural language. To enhance this capability, the AUTOIF Dong et al. (2024) framework introduces two approaches: SFT + Offline DPO and SFT + Iterative Online DPO. For SFT + Offline DPO, after supervised fine-tuning, AUTOIF generates and filters various data scales, creating numerous positive and negative sample pairs for training. The AUTOIF framework employs both offline and iterative online DPO methods, significantly improving a model’s ability to understand and execute natural language instructions.  Yang et al. (2024b) propose an innovative data augmentation technique that simplifies complex instructions by breaking them down into manageable sub-components. These components are then modified and reassembled, preserving the original context and complexity while adding variability essential for training and evaluating the instruction-following capabilities of large language models. Their method can indeed improve the instruction-following capacity of large language models by generating nuanced instruction variants while maintaining the original topic and context, which also allows models to better understand complex instructions. DPO enables LLMs to better finish the decomposed instructions with human preference.
@@
LLMs struggle to generate responses that are aligned with given instructions, especially for complex instructions with multiple restraints. To handle the instruction following needs, Conifer Sun et al. (2024) introduces a new dataset dedicated to improving the instruction-following ability under multiple complex restraints that are closed to the real-world application situations. Conifer constructs a brand-new dataset with GPT-4’s insights on instructions with complex restraints, making it accessible for instruction-following training. The models are further trained by direct preference optimization(DPO). Conifer-7B-DPO is aligned with DPO algorithm, achieving the goal of enhancing instruction following ability of LLMs.
@@
Large language models are prompt-sensitive as the quality of prompts greatly influences the instruction-following ability for LLMs’ performance on downstream tasks. The naive prompts induce the incorrect understanding of instructions. The free-form Free-form Instruction-oriented Prompt Optimization (FIPO; Lu et al. (2024c)) enables models to follow the core instruction within the naive prompts that are considered as model-agnostic manner by accurately optimization. DPO is emplyed in the process of prompt optimization based on the large-scale prompt preference dataset, achieving remarkable instruction-following ability.
@@
With instruction-tuning, large language models learn to follow instructions to finish specific tasks. The ability or performance of LLMs can be evaluated on instruction following and corresponding responses. RS-DPO Khaki et al. (2024) further imroves models’ instruction-following ability by combining rejection sampling and direct preference optimization. Evaluation on instruction-following benchmark demonstrates the effectiveness of thei combination in enhancing models’ performance.
@@ 5.1.3 Application on anti-hallucination
The conventional alignment fails to enhance LLMs’ factual accuracy and leads to the generation of false fact (i.e. hallucination) Lin et al. (2024c). Moreover, former study demonstrated that introducing new knowledge or unfamiliar texts in fine-tuning stage can encourage large language models to hallucinate Gekhman et al. (2024); Lin et al. (2024c), which makes SFT less factual as it trains on human labeled data that may be novel to LLM. Reinforcement learning also encourages LLM to hallucinate as the reward functions prefer longer responses on a diverse set of instructions that contain more details, which tends to stimulate the LLM to yield more false claims.
To tackle these factuality issues, alongside Retrieval-Augmented Generation (RAG; Gao et al. (2023); Zhao et al. (2024a)),  Lin et al. (2024c) propose factuality-aware alignment (FLAME), which incorporates factuality-aware supervised fine-tuning (SFT) and reinforcement learning through direct preference optimization (DPO). DPO is central to this process, ensuring that LLM outputs align with factuality-aware objectives and enhance overall response quality. DPO, combined with factuality-aware alignment, offers a robust framework for improving the quality and factuality of LLM outputs.
@@ 5.1.4 Application on code-generation
Code language models(CLMs) are commonly evaluated by the passing rate to certain task or the accuracy to the generated code task. The running time for CLMs is not a standard criteria for code generation task. Code-Optimise Gee et al. (2024) evaluates the quality of generated by considering both correctness and runtime. Code-Optimise is optimized through direct preference optimization fine-tuning by learning from two dimensional signals including preference pairs over Quick versus Slow as well as Passed versus Failed. An improvement for performance can be observed by DPO fine-tuning.
@@
Stable Code Pinnaparaju et al. (2024) is a serials of based code models that is aimed at code understanding, code interpretation, code completion, mathematical reasoning, and relevant software engineering tasks. Stable Code’s programming capability is further reinforced through SFT and DPO. Training on the preference datasets such as UltraFeedback and HH-Anthropic dataset, Stable Code with 3B parameters outperform large language code models with 7B and 15B parameters, proving the effectiveness of DPO fine-tuning in post-training stage.
@@
Code LLMs usually generate executable code responses to finish the required engineering tasks, assisting human engineers to complete coding domain instructions. However, code LLMs are struggling to generate fast codes for real-world performance because code LLMs only treat code snippets as texts.  Nichols et al. (2024) introduce a new alignment to enhance code LLMs’ performance for generating fast code. Using the variant of DPO fine-tuning, Direct Performance Alignment (DPA; Nichols et al. (2024)) fine-tunes code LLMs by training on faster and slower code paires for preference optimization, achieving new benchmark for code performance.
@@
Large language models have demonstrated comprehensive ability for various downstream tasks including code generation. General LLMs fails to accomplish all code generation tasks especially on competition level programming tasks as these tasks include many corner cases that are not trained in the pre-training stage. Instruct-Code-Llama Liu et al. (2024e) introduces the new fine-tuning approach using Reinforcement Learning with Online Judging Feedback. In RL stage, Instruction-Code-Llama is further aligned using DPO with high-quality data, which achieves less complexity in time and space of generated codes and strengthens the efficiency. The above subsections elaborate the current application of DPO fine-tuning in downstream tasks for forefront LLMs research. Actually, DPO is also employed in evaluation task. RewardBench Lambert et al. (2024) constructs a new benchmark for reward models to evaluate reward models’ capabilities in multiple domains such as chat, reasoning, safety as well as helpfulness and harmlessness. Direct judgement preference optimization Wang et al. (2024e) deploys the DPO fine-tuning to enable the model to learn from the positve and negative samples, enhancing models’ evaluation capabilities across different domain of tasks including reasoning, judgement, and models’ response-comprehension.
@@ 5.2.1 Multi-modal Understanding
Image understanding: A prevalent issue in Large Vision Language Models (LVLMs) is hallucination, where the generated responses fail to align with the provided image and text contexts, significantly restricting the usages of LVLMs Liu et al. (2024f); Yin et al. (2023); Cui et al. (2023b). Generally, humans prefer non-hallucinatory responses over hallucinatory ones. Consequently, some studies have reinterpreted the problem of model hallucinations as a preference optimization task, where DPO is widely employed.
@@
Yu et al. (2024a, b) proposed a token-level variant of DPO to mitigate hallucinations in dense captioning tasks. They found that some key words in the chosen responses required more intensive reinforcement. Specifically, they gathered human preferences  Yu et al. (2024a) or AI feedback  Yu et al. (2024b) in the form of segment-level corrections for hallucinations and implemented a dense variant of DPO.  Xiao et al. (2024) proposed a detect-then-rewrite pipeline for differentiating and mitigating hallucinations, and employed an variant of DPO which adaptively assigns each rejected response a hallucination severity score.
In Llama 3 Herd of Models AI@Meta (2024), authors adopted DPO in post-training stage and have find that instead of always freezing the reference model, updating it in an exponential moving average (EMA) fashion every k-steps helps the model learn more from the data, resulting in better performance in human evaluations. Besides, they observed that the vision DPO model consistently performs better than its SFT starting point in human evaluations for every finetuning iteration.
Video understanding:  Zhang et al. (2024b) introduced a novel framework named LLaVA-Hound that utilizes detailed video captions as a proxy of video content. Specifically, they extracted dense captions on video and then employed a large language model with these captions to score video QA. Their empirical results demonstrate this method’s robust alignment with the reward from GPT-4V OpenAI (2023a), which can take images (i.e. video frames) as input. Additionally,  Zhang et al. (2024d) developed LLaVA-Next-Video by supervised fine-tuning (SFT) LLaVA-Next-Image on video data, achieves better video understanding capabilities compared to LLaVA-Next-Image. Further, they employed LLaVA-Hound Zhang et al. (2024b) which aligns the model response with AI feedback using DPO, showing significant performance boost.
Audio understanding: In Qwen2-Audio’s development Chu et al. (2024), the authors directly employed vanilla DPO to align the SFT model with factuality and human preference.  Ye et al. (2024) developed a MLLM named CAT that responds to audio-visual content, and they have found that although MLLMs are equipped with multimodal understanding, they cannot perceive ambiguity. Therefore, they have reinterpreted ambiguity mitigation as a model preference optimization task and proposed an AI-assisted Ambiguity-aware Direct Preference Optimization (ADPO) method. Specifically, they took ambiguous responses that express the lack of clarity of specific audio-visual objects as negative responses, and then they utilized GPT to rewrite them into positive responses. After the multimodal training, they performed ADPO to align the model to generate the positive response (i.e. the precise description after the rewrite) and reject the negative response (i.e. the ambiguous description).
@@ 5.2.2 Multi-modal Generation
Previous works, such as  Wallace et al. (2023); Yang et al. (2024c), have explored the align T2I generation using DPO. These approaches typically works under the bandit assumption, focusing on a latent reward applied to the entire diffusion denoise process. However, this ignores the sequential nature of the generation process. To address these limitations,  Yang et al. (2024e) propose a fine-grained approach with a dense reward. They develop a tractable objective focusing on the initial stages of the denoising chain. Besides, their method incorporates temporal discounting into implicit reward objectives, which better conform to the hierarchical structure of T2I generation.
Audio generation:  Majumder et al. (2024) leveraged an existing text-to-audio model, Tango, to synthetically create a preference dataset. In this dataset, each prompt is associated with a chosen audio output and several rejected audio outputs, which the diffusion model learns from. The rejected outputs theoretically contain some concepts from the prompt either missing or ordered incorrectly. The authors fine-tuned the text-to-audio model Tango with diffusion-DPO Wallace et al. (2023) on the preference dataset. Their results demonstrate that this approach yields improved audio outputs over Tango, as evidenced by both automatic and manual evaluation metrics.
Current emotional text-to-speech (TTS) models  Diatlova and Shutov (2023); Li et al. (2024b) only learn the correct emotional outputs without fully comprehending other emotion characteristics, limiting their ability to capture the nuances between different emotions.  Gao et al. (2024) propose a controllable Emo-DPO approach, which employs direct preference optimization to differentiate subtle emotional nuances between emotions through optimizing towards preferred emotions over less preferred emotional ones.  Zhang et al. (2024c) have employed DPO in speech generation to align speech generation to human preferences, and they have found that DPO contributes to improvement in speech generation and validated that iterative DPO can consistently enhance the generated speech’s quality.